A shaft of 100mm diameter and 1 meter long has one of its ends fixed and the other end carries a disc of mass 500kg at a radius of gyration of 450 mm. The modulus of rigidity for the shaft material is 80 GN/m2. Determine the frequency of torsional vibration.


1
Expert's answer
2021-11-22T10:31:03-0500

the mass carries at one end and fix at other end hence it is cantilever beam frequency of torsional vibration we know than frequency of torsional vibration is given by


"\u03c9_{nt}=(K_t\/I_p)^{1\/2}"


"K_t=785000N\u2212m"

"I_P=I_X+I_Y=I_Z"


"I_p=2I_X=2\u00d7m\u00d7k^2=2\u00d7500\u00d70.5^2= 250 kg.m^2"

"\u03c9_{nt}=(785000\/250)^{1\/2}"

"=56.03 rad\/s"


"f=56.03/2\times3.14 =8.922 hz"


hence the frequency of torsional vibration


f=8.922 hz